# MoveTouch: Robotic Motion Capturing System with Wearable Tactile Display to Achieve Safe HRI

### v1

@@ 3.3.3 Collaborative Task Procedure
Before the experiment, every participant underwent a training session covering the four patterns and the corresponding movements associated with each pattern.
The robot picks cubes from certain positions on the table and places them in a box near the participant. The participants were asked to touch and familiarize themselves with the box to memorize its position. Throughout the experiment, participants were asked to pick cubes near the box and place them inside the box. The whole experiment was conducted while the user’s eyes were covered.
During collaboration, when the robot’s TCP is moving towards the participant’s hand and gets to a distance closer than 40 cm (the activation area), the system activates a certain haptic pattern on the participant’s wrist to make participants move their hand in response to each pattern as follows: 1) If pattern 1L is rendered, the participant should move their hand to the left. 2) When the pattern is 2L rendered, the participant should move the hand to the right. 3) For the pattern 3L, the movement should be downward. 4) For pattern 5H, the participant should move the hand backward.
For safety concerns, the robot is forced to stop when its TCP comes within a critical threshold (25 cm) from the participant’s hand and continues to move when the user’s hand gets outside the critical area.
Figure 4: The distance between the participant and the robot’s TCP. The red line is the critical distance, while the orange line represents the haptic activation distance. The four highlighted areas are the areas where the robot is approaching the human.
We measured the response time that the participants needed to get their hands out of the dangerous area. The time was measured between the moment the MoveTouch gives the haptic pattern and the moment users start to move their hands. The measured time for all users for each pattern can be seen in Fig. 5. We can see that the mean response times were 0.24 sec, 0.61 sec, 0.85 sec, and 2.41 sec for the first, second, third, and fourth patterns, respectively. We can see that the response time of the fourth pattern is the highest among all the patterns.
Figure 5: The response time for participants to remove their hands for the four patterns.
@@ 3.3.4 Results
The distance between the participant’s hand and the robot’s TCP is illustrated in Fig. 4. We can observe that the critical distance was preserved throughout the experiment, which indicates safer collaboration, taking into account that the user was doing the task while eye-covered. All participants were able to respond to the haptic patterns correctly, with the exception of one, where the pattern 5H was used to move downward instead of backward. We can see the effective response to the haptic patterns in the highlighted areas in Fig. 4. Those areas correspond to the moment in which the robot is approaching the participant. For example, in the blue highlighted area, the participant is trying to continue the task and return the hand near the box many times, but the robot is still approaching, triggering the haptic pattern multiple times. Consequently, the participant moves their hand again, resulting in this oscillating robot-user distance.
@@ 4 Conclusion and Future Work
This research presents an innovative wearable robotic system featuring both motion capture and a haptic guidance system designed for human-robot interaction. The system includes a wearable 2-DoFs robot responsible for adjusting a marker, ensuring its constant visibility to the camera. The haptic guidance system incorporates five vibration motors positioned at the user’s wrist. A total of ten haptic patterns were selected for experimentation on both the volar and dorsal parts of the wrist. A user study was conducted to evaluate the recognition and discriminability of these patterns. Patterns with the highest recognition rates were subsequently chosen. Another experiment was made to evaluate the integration of the system into a real-world collaborative task. We saw that the critical distance had not been violated during the whole experiment and that the users were efficiently responding to the haptic system.
For the future, we aim to further study the response time of the users for all four patterns and to study how a different application and the speed of the robot might affect the result.

### v2

@@ 3.3.3 Collaborative Task Procedure
Before the experiment starts, each participant underwent a training session covering the four patterns and the corresponding movements associated with each pattern.
The robot picked cubes from certain positions on the table and placed them in a box near the participant. The participants were asked initially to touch and familiarize themselves with the box to memorize its position. Throughout the experiment, participants were asked to pick cubes near the box and place them inside the box. The whole experiment was conducted while the user’s eyes were covered.
During collaboration, when the robot’s TCP is moving towards the participant’s hand and gets to a distance closer than 40 cm (the activation area), the system activates a certain haptic pattern on the participant’s wrist to make participants move their hand in response to each pattern as follows: 1) If pattern 1L is rendered, the participant should move their hand to the left. 2) When the pattern 2L is rendered, the participant should move the hand to the right. 3) For the pattern 3L, the movement should be downward. 4) For pattern 5H, the participant should move their hand backward.
For safety concerns, the robot was programmed to halt when its TCP came within a critical threshold (25 cm) from the participant’s hand and continue when the participant’s hand moved outside the critical area.
Figure 5: The distance between the participant and the robot’s TCP during one experiment. The red line is the critical distance, while the orange line represents the haptic activation distance. The four highlighted areas are the areas where the haptic patterns were activated.
@@ 3.3.4 Results
The distance between the participant’s hand and the robot’s TCP is illustrated in Fig. 5. We can observe that the critical distance was preserved throughout the experiment, which indicates safer collaboration, taking into account that the user was doing the task while eye-covered. All participants successfully responded to the haptic patterns, except for one instance where the participant responded to the pattern 5H by moving his hand downward instead of backward. We can see the effective response to the haptic patterns in the highlighted areas in Fig. 5. These highlighted areas correspond to moments when the robot approached the participant. For example, in the blue highlighted area, the participant attempted to continue the task and return his hand near the box many times, but the robot was still approaching, triggering the haptic pattern multiple times. Consequently, the participant moves their hand again, resulting in an oscillating distance between the robot and the user.
We measured the response time that the participants needed to get their hands out of the dangerous area. The time was measured between the moment the MoveTouch gives the haptic pattern and the moment users start to move their hands. The measured time for all users for each pattern can be seen in Fig. 6. We can see that the mean response times were 0.24 seconds, 0.61 seconds, 0.85 seconds, and 2.41 seconds for the first, second, fourth, and third patterns, respectively. We can see that the response time of the fourth pattern is the highest among all the patterns.
Figure 6: The response time for participants to remove their hands for the four patterns during the collaborative task.
3.3.5 Conclusion
The critical distance between the eye-covered participant and the robot’s TCP was preserved throughout the collaboration, indicating the effectiveness of the designed vibrotactile patterns in guiding user actions when rendered on the wrist. The patterns remained discernible even when the participant’s attention was focused on performing another task. The measured perception time of each pattern can serve as a design parameter to establish an upper limit for the robot’s speed during collaboration, to assure that the robot does not critically approach the user, and thus, enhance user safety during collaborative tasks.
@@ 4 Conclusion and Future Work
This research presents an innovative wearable robotic system featuring both motion capture and a haptic guidance system designed for human-robot interaction. The system includes a wearable 2-DoF robot responsible for adjusting a marker held at its end-effector, ensuring its constant visibility to the camera. The haptic guidance system incorporates five vibration motors fixed on the wearable robot wrist rubber band to render vibrotactile patterns at the user’s wrist to guide their hand movement.
A total of ten haptic patterns were selected for experimentation on both the volar and dorsal parts of the wrist. A user study was conducted to evaluate the recognition and discriminability of these patterns. The average recognition rate of all patterns on the dorsal (up) part of the wrist was 66.6%, while on the volar (down) side of the wrist, it was 75.64%. In addition, almost all the vibration patterns were better recognized on the volar part of the wrist. Consequently, the decision was made to fix the vibration motors at the volar part of the wrist. Four patterns with high recognition rates were selected for further study.
Another experiment was made to evaluate the integration of the system into a real-world collaborative task incorporating the four selected patterns. We report that the critical distance had not been violated during the whole experiment and that the users were efficiently responding to the vibrotactile patterns. We also measured the perception time of each pattern and it was in the range 0.24-2.41 seconds.
For future work, we intend to expand the testing of the Movetouch robot across a broader user base and in diverse tasks to better evaluate the effectiveness of the proposed guidance system. We also aim to study how to enhance the recognition rate of the patterns by studying the effect of the pattern frequency. Additionally, we seek to minimize user response times to suggested patterns, considering that some patterns are perceived much faster than others.
Acknowledgements
The research reported in this publication was financially supported by the Russian Science Foundation grant No. 24-41-02039.